OpenStudy (anonymous):

Prove for each n ∈ ℤ+, n > 3, 2^n >= n^2

OpenStudy (anonymous):

\[2^n \ge n^2\] then: \[n \ge 2\log_2n\] then: \[\frac{n}{2} \ge \log_2n\]

OpenStudy (anonymous):

this is true for 4

OpenStudy (anonymous):

\[\frac{d(n/2)}{dn}=(1/2)\]

OpenStudy (anonymous):

\[\frac{d(\log_2n)}{dn}=\frac{1}{n*\ln2}\]

OpenStudy (anonymous):

therefore, for n > 2/ln(2) (2.88) n/2 grows faster than log_2n

OpenStudy (anonymous):

So basically, it satisfies this relationship for 4, and then the LHS grows faster than the RHS

OpenStudy (anonymous):

Now, first lets Prove n^2>2n+1 for n>3 WHEN n=4, n^2>2n+1 ( TRUE) Now, when n=n+1 (n+1)^2>2(n+1)+1 n^2+2n+1>2n+2+1 n^2>2 (WHICH IS ALWAYS TRUE when n>3) Thus, n^2>2n+1 for n>3 NOW, WHEN n=4 then 2^n>=n^2 (TRUE) WHEN n=5 then 2^n>n^2 (TRUE) WHEN n=n+1 then 2^(n+1)>(n+1)^2 2*2^n>n^2 + (2n+1) This is true because when n=5, 2^n>n^2 AND 2^n>n^2>2n+1 Thus 2*2^(n+1)>(n+1)^2 is true for all n>4 SO, 2^n>=n^2 for n>3